OpenStudy (anonymous):

y=3(x-6) +3 if you add 3 to both sides which makes a line with slope 3 y-intercept at 3 x-intercept at 6 What would be an example of a line parallel to your answer up above?

OpenStudy (anonymous):

a line is parallel if it has the same slope, so then since the slope is 3 you can change the way it intercepts the x or y axis by changing 6 or 3 so an example like this can be y=3(x-6) +4 y=3(x-6) +3 + a where a is any constant also, applying this to x you can change y=3(x-6+ a) +3